OpenStudy (anonymous):

Write the function in vertex form and identify its vertex. g(x) = x2 - 10x + 11

OpenStudy (anonymous):

same idea as the last one. lets see if i can do it without making a mistake

OpenStudy (anonymous):

\[-\frac{b}{2a}=\frac{10}{2}=5\] and \[g(5)=5^2-10\times 5 +11=25-50+11=-25+11=-14\] so vertex is \[(5,-14)\] and vertex form will be \[g(x)=(x-5)^2-14\]

OpenStudy (anonymous):

you can check that this is correct by multiplying out \[(x-5)^2-14\] and see that you get \[x^2-10x+11\]

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

always works like this. you can also "complete the square" if you like, but it is unnecessary